OpenStudy (anonymous):

Analyze the sequence below and answer the question that follows. 1, 6, 22, 69, 210 Which element in the sequence is a counterexample to the rule? Rule: Add 1, then multiply by 3 to get the next element. A. 6 B. 22 C. 69 D. 210

OpenStudy (anonymous):

So let's just solve each one until we find one that isn't true :)

OpenStudy (anonymous):

So let's start with 1 1 + 1 = 2 2 * 3 = 6 The next element is 6 so 1 is not the counterexample

OpenStudy (anonymous):

6 + 1 = 7 7 * 3 = 21 However, the next number in the list is 22 and we got 21... So 22 is the counterexample your answer is B

OpenStudy (anonymous):

thank you again :)